OpenStudy (anonymous):

Evaluate x^(1) - x^(-1) + x^(0) for x = 2. will someone please explain this

OpenStudy (anonymous):

\(2^{-1} = 1/2\)

OpenStudy (anonymous):

\(2^0=1\)

OpenStudy (anonymous):

\(2^1=2\)

OpenStudy (anonymous):

 that's what the choices I have